OpenStudy (anonymous):

What is a possible first step in eliminating a variable in the following system of equations? x + 7y = –15 2x + 7y = –9

OpenStudy (anonymous):

MY answer choices Multiply the first equation by –1. Multiply the second equation by –1 Multiply the first equation by –2. All of the above

OpenStudy (anonymous):

Try doing each one... first, do one of your options, like multiply the first equation by -1. Now add it to the 2nd equation and see if a variable disappears. So for the first case, you have \[(x+7y=-15) *(-1)\] which gives you \[-x-7y=15\] If you add that to the 2nd equation, you'll eliminate the 7y. Try for the other 2.

OpenStudy (texaschic101):

all of the above :)